OpenStudy (anonymous):

I do not understand the answers I was given to this question. The average cost of tuition and fees at private four-year colleges was $16,200, and in 2005 it was $20,100. Sketch a line that passes through the points (2000, 16200) and (2005, 20100). My question is - find the slope-intercept form of the line in the sketch. What is the y-intercept and does it have meaning in this situation?

OpenStudy (anonymous):

Alright, I'll solve it showing the steps, and see if that makes a bit more sense. The slope intercept equation for a line is this: \[y = mx + b\] The m is the slope of the line. You find the slope using this: \[\Delta Y \div \Delta X = Y _{1} - Y _{2} div X _{1} - X _{2}\] Using your values and substituting them in, you have: \[(20100 - 16200) \div (2005 - 2000) = (3900) \div (5) = 780.\] So, the slope is 780. Then we need to find the y-intercept, which in the equation is represented by the variable b. We have a coordinate in the (x,y) format, so we will use those to solve for b: \[y = mx + b\] \[16200 = (780)(2000) + b\] \[16200-1560000 = b\] \[b = -1543800\] Obviously b is kinda a wacky number, but it is the point at which the line reaches the y axis.